OpenStudy (chrispaul03):

FAN AND MEDAL -7+4n=9

OpenStudy (chrispaul03):

@AloneS

alones (alones):

So simplify both sides of the equation 4n−7=9 Then add 7 to both sides. 4n−7+7=9+7 4n=16 Step 3: Divide both sides by 4. 4n/4=16/4

OpenStudy (gnorris):

-7 + 4n = 9 Solving for variable 'n'. Move all terms containing n to the left, all other terms to the right. Add '7' to each side of the equation. -7 + 7 + 4n = 9 + 7 Combine like terms: -7 + 7 = 0 0 + 4n = 9 + 7 4n = 9 + 7 Combine like terms: 9 + 7 = 16 4n = 16 Divide each side by '4'. n = 4